Find the partial derivative of f(x,y,z)=x^2y+6z^3xy
1
Expert's answer
2020-10-26T19:41:51-0400

Solution:Partial derivatives:

"\frac{\partial \:}{\partial \:x}\left(x^{2y}+6z^{3xy}\right)=2yx^{2y-1}+18yz^{3yx}\ln \left(z\right)"


"\frac{\partial \:}{\partial \:y}\left(x^{2y}+6z^{3xy}\right)=2x^{2y}\ln \left(x\right)+18xz^{3xy}\ln \left(z\right)"


"\frac{\partial \:}{\partial \:z}\left(x^{2y}+6z^{3xy}\right)=18xyz^{3xy-1}"